OpenStudy (anonymous):

Perform the indicated operations. Simplify, and state the restrictions on the variable 13/3x+12 + 2x/x+4

OpenStudy (anonymous):

i will give medals

OpenStudy (mathteacher1729):

What will these medals buy me in real life? :-p

OpenStudy (anonymous):

lol idk

OpenStudy (wasiqss):

(13+6x)/(3(x+4)

OpenStudy (hoblos):

haha lol to add fractions you must have same denominator 3x+12 = 3(x+4) so to obtain same denominator multiply the second expression by 3/3

OpenStudy (hoblos):

13/3x+12 + 2x/x+4 = 13/3(x+4) + 6x/3(x+4) = (13+6x)/3(x+4)

OpenStudy (hoblos):

with a restriction at x=-4